A man bought some tools for a total amount of P90,000. If he had bought 50 more tools, it would have cost him P20 less per piece. How many tools did he buy?
Answer by stanbon(75887)

A man bought some tools for a total amount of P90,000. If he had bought 50 more tools, it would have cost him P20 less per piece. How many tools did he buy?
-------------
Let the original number of tools he bought be "x".
---
Average cost per tool = 90,000/x
----------------------
Average cost if he had bought 50 more would be
90,000/(x+50)
----
Equation:
[(90000/x)-(90000/(x+50)] = 20
---
90000*50 = 20x(x+50)
----------
x^2+50x -225000 = 0
---
x = 450 (original # of tools purchased)
=============================================
Cheers,
Stan H.
===========